Simon Greenwold

Newton, MA USA

Homespun (2005)

Dollhouse, electric motor, camera, computer, monitors

Homespun is dedicated to anyone who has ever felt turned upside down. The dollhouse rotates end-over-end on an axle. In the video, the rotation is removed, so the furniture prowls slowly around the room. It is about the heaviness and danger of the objects with which we surround ourselves and the fragility of the shells we live in.

Eric Gunther, Justin Manor, John Rothenberg

Cambridge, MA USA

www.sosolimited.com

Algorithmic Materialization Study #0 (2005)

televisions, software, sound

After establishing a set of algorithmic and architectural rules, each member of sosolimited was given a graphical primitive - points, lines, or planes - with which to compose in graphical and sonic space. The sounds drive the motion graphics and the graphics modulate the sounds.

Steve Hollinger

Boston, MA USA

www.stevehollinger.com

Downtown (2005)

Mixed media

“Downtown” is one in a series of Steve Hollinger’s sculptures that use strands of thread to examine the most delicate of forms, structures and connections.

jackbackrack

Cambridge, MA USA

Full Backup (2004)

Video Camera, Computer, Projector

fullbackUP is an interactive video installation that responds to people’s movements by amplifying and replicating them. Internal video feedback loops generate direct and indirect patterns, introduce delays and effects to produces a visual landscape that is revealing and constantly changing.

Brian Knep

Boston, MA USA

www.blep.com

Big Smile (2003)

Computer, Video Projector, Video Camera, Custom Software

This face smiles only when no one is looking directly at it. By doing so, it shows its discomfort with being the center of attention and even its disdain with the audience. It asks, “Why are you looking at me,” when of course without viewers the art piece wouldn’t exist.

Jeff Lieberman

Cambridge, MA USA

http://bea.st

Slink (2004)

Aluminum, Corroded Steel, Acrylic, Electronics, Custom Voice Coil with Flexure Mounts, 1980 LEDs, Extension Spring

Slink is an experiment with matched mechanical, electrical, and visual resonances, using a spring and light. This is an actual vibrating spring, not a computer simulation.

Kevin McCormick

Cambridge, MA USA

www.hydrochloride.net/rhombus/

Corona (2003)

naked electronics, LEDs

Corona is a geodesic sphere constructed entirely of printed circuit boards and light-emitting diodes. Serving as both a lamp and a three-dimensional display, its 180 LED clusters can take on and project any color of the rainbow under real-time computer control. Corona throws off the bonds of the hot and bulky incandescent bulb, and imagines what our lighting fixtures might be like in the networked, semiconducting future.

Andrew Neumann

Boston, MA USA

www.bitforms.com/artist neumann.html

“Industrial Tree” (2005)

Plywood, fans, analog switch, wire, power supply

Dan Roe

Cambridge, MA USA

www.danroe.net

Captivity: Specimens one and two (2004)

Steel, solar engine (motors, glass solar panels, circuitry)

These specimens are artificial life forms created from wire and circuitry. Consider either one, and let us wonder at the most life-like aspect of this individual ”artificial life form.” Is it some physical characteristic, the shape of the wings, the tail, the placement of motors roughly where feet would be, or perhaps the infrared photosensors in the place of eyes with brain-like control circuitry behind? Or is the most life-like characteristic behavioral, and to be observed in the way the body moves and flexes against its chains, while simultaneously seeking a light source to power its movement? Neither proposition addresses the situation in which this sculpture finds

William Tremblay, Rob Gonsalves

Allston, MA USA

Wave Puppet (2005)

Acrylic, aluminum, steel, EPDM rubber, servo motors, computer, custom software

Translating powerful physical forces to an anthropomorphically comprehensible and safely inanimate form, Wave Puppet is a marionette of the ocean’s surface directed by the math that underlies all waves.